OpenStudy (anonymous):

One number is 3 times another number and their sum is 8. Find the numbers.

OpenStudy (anonymous):

Is this right? x = smaller number 3x = larger number n + 3n = 8 4n = 8 4n/4 = 8/4 n = 2 3n = 6

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

Thanks :D